OpenStudy (anonymous):

find the equation of the line in slope intercept form, that passes through (4,-1) and (6,-2) can someone help me please :)

myininaya (myininaya):

Do you know what slope intercept form is?

OpenStudy (anonymous):

yes, when i did it i got \[y= -\frac{ 1 }{ 2 }x+\frac{ 7 }{ 2 }\]

myininaya (myininaya):

The slope is right...Checking the y-intercept now...

OpenStudy (anonymous):

ok

myininaya (myininaya):

You should retry to find the y-intercept. It is incorrect. Start here: y=mx+b where m=-1/2 y=-1/2 * x +b Use one of the points gven to plug in for x and to plug in for y.

myininaya (myininaya):

Either use (4,-1) or (6,-2) but don't mix and match

OpenStudy (anonymous):

i got 1

myininaya (myininaya):

yes b=1

OpenStudy (anonymous):

so if i were to show my work on a homework assignment i would find the slope and then put it in point slope form? or just plug it back into the equation in slope intercept

myininaya (myininaya):

Either way.

OpenStudy (anonymous):

because when i did it in point slope form like my teacher said i did not get one i got 7/2

myininaya (myininaya):

Then you did it incorrectly. You found the slope m=-1/2 Plug it in y-y1=m(x-x1) Don't forget to distribute y-y1=mx-mx1

myininaya (myininaya):

add y1 on both sides y=mx-mx1+y1

OpenStudy (anonymous):

yeah i got it now. and one more question standard form is ax+by=c right?

myininaya (myininaya):

Yes.

OpenStudy (anonymous):

okay because my book said ax+by+c=0 and my teacher said ax+by=c

myininaya (myininaya):

Go with what your teacher said. Either one is fine. Both of those forms don't differ a whole a lot. All my math life, I have prefered the slope-intercept form.

myininaya (myininaya):

Your teacher will be grading you and it is their preference in what form they want the answer.

OpenStudy (anonymous):

okay thank you so much :)